Assume that x has a normal distribution with the specified mean and standard deviation. Find the indicated probability. (Round your answer to four decimal places.)

μ = 15.7; σ = 4.0

P(10 ≤ x ≤ 26) =
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 15.7
std deviation   =σ= 4.000
probability =P(10<X<26)=P((10-15.7)/4)<Z<(26-15.7)/4)=P(-1.43<Z<2.58)=0.9951-0.0764=0.9187

(Note:

if using ti-84 use command :normalcdf(10,26,15.7,4)
if using excel use command :norm.dist(26,15.7,4,true)-norm.dist(10,15.7,4,true) )